OpenStudy (anonymous):

Question :Every weekend , Morga buys intresting clothes at her local thrift store and then resells them on an auction website. If she brings $150.00 and spends s,write an expression for how for how much change she has. Evaluate your expression for s=27.13 and s=55.14

OpenStudy (anonymous):

Equation = Money she brings - money spent = change; so, $150 - s = change; s = 27.13, 150-27.13=122.87; s=55.14, 150-55.14=94.86... Hope this helps...